Find the eqn of the plane passing through the point 2i-2j+2k and all parallel to the plane r.(2i-4j+2k)=5

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear LINTA

equation of plane parallel to r.(2i-4j+2k)=5  is

  r.(2i-4j+2k)=d

this plane passes through  2i-2j+2k so

 (2i-2j+2k).(2i-4j+2k)=d

d = 4+8+4 =16

so desired plane r.(2i-4j+2k)=16

   
Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin